# A deep backward regression-based scheme for high-dimensional nonlinear partial differential equations

### 16 Mar 2026 v1

@@ 4.2 Convergence analysis
$\quad+2L^{2}h^{2}(1+\frac{1}{\gamma h})\mathbb{E}\left[|\overline{Y}_{i}^{\pi,m}-\overline{\mathcal{U}}^{m}_{t_{i}}|^{2}+|Z_{i}^{\pi,m}-\widehat{\mathcal{V}}^{m}_{t_{i}}|^{2}\right].$ (4.8)
@@
$\mathbb{E}\left[\left|Z_{i}^{\pi,m}-\widehat{\mathcal{V}}^{m}_{t_{i}}\right|^{2}\right]\leq$ $\frac{2d}{h}\mathbb{E}\left[\left|Y_{i+1}^{\pi,m}-\widehat{\mathcal{NN}}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})\right|^{2}\right]$ (4.9)
$-\frac{2d}{h}\mathbb{E}\left[\left|\mathbb{E}_{i}\left[Y_{i+1}^{\pi,m}-\widehat{\mathcal{NN}}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})\right]\right|^{2}\right].$
Plugging (4.9) into (4.8) and choosing $\gamma=4dL^{2}$, we have for small enough $h$
$\mathbb{E}\left[\left|\overline{Y}_{i}^{\pi,m}-\overline{\mathcal{U}}^{m}_{t_{i}}\right|^{2}\right]\leq$ $(1+Ch)\mathbb{E}\left[\left|Y_{i+1}^{\pi,m}-\widehat{\mathcal{NN}}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})\right|^{2}\right].$ (4.10)
Next, by Young inequality as the Step 1 in the proof of Theorem 3.1, we derive
$\mathbb{E}\left[\left|\overline{Y}_{i}^{\pi,m}-\widehat{\mathcal{Y}}_{i}\right|^{2}\right]$ $\leq(1+Ch)\mathbb{E}\left[\left|Y_{i+1}^{\pi,m}-\widehat{\mathcal{NN}}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})\right|^{2}\right]$ (4.11)
$\quad+CN\mathbb{E}\left[\left|\widehat{\mathcal{Y}}_{i}-\overline{\mathcal{U}}^{m}_{t_{i}}\right|^{2}\right].$
@@
By the same arguments as in Step 2 in the proof of Theorem 3.1, it follows that for $h$ small enough
@@
$\mathbb{E}\left[\left|\widehat{\mathcal{Y}}_{i}-\overline{\mathcal{U}}^{m}_{t_{i}}\right|^{2}\right]$ $\leq C\mathbb{E}\left[N\left|\widehat{R}_{i+1,2}^{y}\right|^{2}+\left|\widehat{R}_{i,2}^{z}\right|^{2}+N\left|\widehat{R}_{i+1,1}^{y}\right|^{2}+\left|\widehat{R}_{i,1}^{z}\right|^{2}\right].$ (4.12)
Plugging (4.12) into (4.11) and recalling that $\widehat{\mathcal{U}}_{t_{i}}^{m}=\max\left\{\overline{\mathcal{U}}_{t_{i}}^{m},g\left(X_{i}^{\pi,m}\right)\right\}$, $Y_{i}^{\pi,m}=\max\left\{\overline{Y}^{\pi,m}_{i},g\left(X_{i}^{\pi,m}\right)\right\}$ and $\left|\max(a,c)-\max(b,c)\right|\leq\left|a-b\right|$, we have
@@
$\max\limits_{0\leq i\leq N-1}\mathbb{E}\left[\left|Y_{i}^{\pi,m}-\widehat{\mathcal{Y}}_{i}\right|^{2}\right]\leq C\sum\limits_{i=0}^{N-1}\mathbb{E}\left[N\left|\widehat{R}_{i+1,2}^{y}\right|^{2}+\left|\widehat{R}_{i,2}^{z}\right|^{2}+N\left|\widehat{R}_{i+1,1}^{y}\right|^{2}+\left|\widehat{R}_{i,1}^{z}\right|^{2}\right].$ (4.13)
Combining (4.13) with (4.6), we deduce
$\max\limits_{0\leq i\leq N-1}\mathbb{E}\left[\left|Y_{t_{i}}-\widehat{\mathcal{Y}}_{i}\right|^{2}\right]\leq C\sum\limits_{i=0}^{N-1}\mathbb{E}\left[N\left|\widehat{R}_{i+1,2}^{y}\right|^{2}+\left|\widehat{R}_{i,2}^{z}\right|^{2}+N\left|\widehat{R}_{i+1,1}^{y}\right|^{2}+\left|\widehat{R}_{i,1}^{z}\right|^{2}\right].$ (4.14)
This establishes the error estimate for the $Y$-component. The error estimate for the $Z$-component can be derived using the same arguments as in Steps 3 and 4 of the proof of Theorem 3.1; thus, the details are omitted here.
@@ 5 Numerical experiments
In this section, we demonstrate the performance of Algorithm 1 to solve the high dimensional nonlinear PDEs (1.1) via two numerical examples.
@@
$\begin{cases}\displaystyle\frac{\partial u}{\partial t}(t,x)+\frac{d^{2}}{2}\,\Delta_{x}u(t,x)+\Big(u(t,x)-\frac{d+2}{2d}\Big)\Big(d\sum_{\ell=1}^{d}\frac{\partial u}{\partial x_{\ell}}(t,x)\Big)=0,&(t,x)\in[0,T)\times\mathbb{R}^{d},\\[10.00002pt] \displaystyle u(T,x)=\frac{\exp\!\left(T+\tfrac{1}{d}\sum_{\ell=1}^{d}x_{\ell}\right)}{1+\exp\!\left(T+\tfrac{1}{d}\sum_{\ell=1}^{d}x_{\ell}\right)}.\end{cases}$ (5.1)
@@
In the numerical experiments we take the terminal time $T=1$, the number of time steps $N=10$, the Monte Carlo samples $M=10000$, and the initial state $x_{0}=(0,0,\dots,0)^{\top}\in\mathbb{R}^{d}$. Thus, The true solution of (5.1) at the initial point is $u(0,x_{0})=0.5$. The numerical solution at the initial point is denoted by $\mathfrak{u}(0,x_{0})$. The numerical solver is implemented using a deep learning framework where two sub-networks are trained at each time step to approximate the solution $u$ and its gradient $\nabla u$. Each network consists of two hidden layers with $d+110$ units and $\tanh$ activations. We utilize the Adam optimizer with a learning rate of $5\times 10^{-4}$ and conduct 6,000 training iterations per step to ensure the convergence of the loss function.
$u(t,x)$ and its estimate at time $t=0.2$.
$u(t,x)$ and its estimate at time $t=0.4$.
$u(t,x)$ and its estimate at time $t=0.6$.
$u(t,x)$ and its estimate at time $t=0.8$.
Figure 1: Estimated solution $\mathfrak{u}(t,x)$ obtained by DBR versus exact solution $u(t,x)$ for Example 1 with $d=1$.
@@
Scheme $d$ $u(0,x_{0})$ $\overline{\mathfrak{u}}(0,x_{0})$ $\operatorname{Std.dev.}$ $\mathbb{E}[|u(0,x_{0})-\mathfrak{u}(0,x_{0})|]$ $\operatorname{Rel.err.}$
@@
Table 1 complements these visual findings by providing a quantitative comparison between the proposed DBR algorithm and the existing DBDP1 method in [28] across multiple high dimensions ($50,80,100$ and $200$). For each scheme and dimension, we report the sample mean $\overline{\mathfrak{u}}(0,x_{0})$ over 10 independent runs, the empirical standard deviation $\operatorname{Std.dev.}(u^{\pi}(0,x_{0}))$, the mean absolute error $\mathbb{E}[|u(0,x_{0})-\mathfrak{u}(0,x_{0})|]$, and the relative error $\operatorname{Rel.err.}(\mathfrak{u}(0,x_{0}))=|u(0,x_{0})-\overline{\mathfrak{u}}(0,x_{0})|/|u(0,x_{0})|$. Regarding accuracy, the DBR algorithm maintains a relative error consistently below $4\%$ across all tested dimensions, significantly outperforming the DBDP1 method which exhibits errors as high as $13.9\%$ at dimension $50$ and $9.7\%$ at dimension $80$. More importantly, the mean absolute error, which represents the average magnitude of the absolute deviation between the true solution and the numerical approximation, is consistently smaller for the DBR method across all dimensions. For instance, at dimension $100$, DBR achieves a mean absolute error of $0.008006$ compared to DBDP1’s $0.045578$, indicating that the DBR algorithm produces approximations that are substantially closer to the true solution on average. This aligns with the papers’ claim that the DBR method produces more accurate estimations, particularly as the dimension increases. Regarding stability, the standard deviations reported for both methods are generally comparable in magnitude across the tested dimensions. The numerical results validate the core theoretical advantages of the DBR method.
Example 2. (see [17]) Consider the following high-dimensional nonlinear PDEs which are equipped with unbounded solution and more complex structure of the form
@@
with $\hat{k}(t,x)=\partial_{t}u+\frac{1}{2d}\operatorname{Tr}\left(D_{x}^{2}u\right)+\frac{u}{\sqrt{d}}D_{x_{i}}u+\frac{u^{2}}{2}$. Thus, the solution of PDE (5.2) is given by $u(t,x)=\frac{T-t}{d}\sum_{i=1}^{d}\left(\sin\left(x_{i}\right)1_{x_{i}<0}+x_{i}1_{x_{i}\geq 0}\right)+\cos\left(\sum_{i=1}^{d}ix_{i}\right).$
Example 2 presents a more challenging class of high-dimensional nonlinear PDEs characterized by an unbounded solution and a complex structural form. To comprehensively evaluate the performance of the proposed DBR algorithm, we conduct systematic numerical experiments across dimensions $d=1,2,8,15,20$ with varying time steps $N$, where the initial point is fixed as $x_{0}=0.5\mathbf{1}_{d}$ for all cases. The neural network architecture consists of two hidden layers with hyperbolic tangent activation functions, where each hidden layer contains $d+10$ units. The model is trained using a batch size of 400 and a total of 3000 iterations.
$u(t,x)$ and its estimate at time $t=0.3$.
$u(t,x)$ and its estimate at time $t=0.5$.
$u(t,x)$ and its estimate at time $t=0.7$.
$u(t,x)$ and its estimate at time $t=0.9$.
Figure 2: Estimated solution $\overline{\mathfrak{u}}(t,x)$ obtained by DBR versus exact solution $u(t,x)$ for Example 2 with $d=1$.
@@
Scheme NN $u(0,x_{0})$ $\overline{\mathfrak{u}}(0,x_{0})$ $\operatorname{Std.dev.}$ $\mathbb{E}[|u(0,x_{0})-\mathfrak{u}(0,x_{0})|]$ $\operatorname{Rel.err.}$
@@
Scheme NN $u(0,x_{0})$ $\overline{\mathfrak{u}}(0,x_{0})$ $\operatorname{Std.dev.}$ $\mathbb{E}[|u(0,x_{0})-\mathfrak{u}(0,x_{0})|]$ $\operatorname{Rel.err.}$
@@
Scheme NN $u(0,x_{0})$ $\overline{\mathfrak{u}}(0,x_{0})$ $\operatorname{Std.dev.}$ $\mathbb{E}[|u(0,x_{0})-\mathfrak{u}(0,x_{0})|]$ $\operatorname{Rel.err.}$
@@
Scheme NN $u(0,x_{0})$ $\overline{\mathfrak{u}}(0,x_{0})$ $\operatorname{Std.dev.}$ $\mathbb{E}[|u(0,x_{0})-\mathfrak{u}(0,x_{0})|]$ $\operatorname{Rel.err.}$
@@
Scheme NN $u(0,x_{0})$ $\overline{\mathfrak{u}}(0,x_{0})$ $\operatorname{Std.dev.}$ $\mathbb{E}[|u(0,x_{0})-\mathfrak{u}(0,x_{0})|]$ $\operatorname{Rel.err.}$

### 21 May 2026 v3

@@ 4.2 Convergence analysis
$\overline{\varepsilon}_{i}^{y}=$ $\ \inf_{\widehat{\mathcal{NN}}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}}\mathbb{E}\left[\left|\widetilde{\mathcal{U}}_{i}^{m}-\widehat{\mathcal{NN}}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{y,i}^{r_{j}})\right|^{2}\right],$ (107)
$\overline{\varepsilon}_{i}^{z}=$ $\ \inf_{\widehat{\mathcal{NN}}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}}\mathbb{E}\left[\left|\widetilde{\widehat{\mathcal{V}}}_{i}^{m}-\widehat{\mathcal{NN}}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{z,i}^{r_{j}})\right|^{2}\right].$
The result is obtained under one of the following additional assumptions.
Assumption 1. The terminal function $g$ is Lipschitz and satisfies $g\geq\Phi$. Moreover, $\Phi$ is $C^{1}$, and $\Phi,D_{x}\Phi$ are Lipschitz.
Assumption 2. The terminal function $g$ is Lipschitz and satisfies $g\geq\Phi$. Moreover, $\sigma$ is $C^{1}$, with $\sigma,D_{x}\sigma$ both Lipschitz, and $\Phi$ is $C^{2}$, with $\Phi,D_{x}\Phi,D_{x}^{2}\Phi$ all Lipschitz.
For later use, we separate the time-discretization errors of the reflected BSDE into the $Y$-component and the $Z$-component:
$\varepsilon_{Y}(h)=\max_{0\leq i\leq N-1}\mathbb{E}\left[\left|Y_{t_{i}}-Y_{i}^{\pi,m}\right|^{2}\right],\qquad\varepsilon_{Z}(h)=\mathbb{E}\left[\sum_{i=0}^{N-1}\int_{t_{i}}^{t_{i+1}}\left|Z_{t}-Z_{i}^{\pi,m}\right|^{2}dt\right].$ (108)
According to [3], we have
$\varepsilon_{Y}(h)=O(h^{1/2})\quad\text{under Assumption 1},\qquad\varepsilon_{Y}(h)=O(h)\quad\text{under Assumption 2},$
while
$\varepsilon_{Z}(h)=O(h^{1/2}).$
Theorem 4.
Let the assumptions (i)–(iv) hold together with either Assumption 1 or Assumption 2. Let $(Y_{t_{i}},Z_{t_{i}})$ and $(\widehat{\mathcal{Y}}_{i},\widehat{\mathcal{Z}}_{i})$ be the solution of the RFBSDE (96) and the solution of the RDBR method (99), respectively. Then, for small enough $h$, we have
$\max_{0\leq i\leq N-1}\mathbb{E}\left[\left|Y_{t_{i}}-\widehat{\mathcal{Y}}_{i}\right|^{2}\right]+\mathbb{E}\left[\sum_{i=0}^{N-1}\int_{t_{i}}^{t_{i+1}}\left|Z_{s}-\widehat{\mathcal{Z}}_{i}\right|^{2}ds\right]$ (109)
$\leq C\varepsilon_{Y}(h)+C\varepsilon_{Z}(h)+C\sum_{i=0}^{N-1}\left(N\widehat{\varepsilon}_{i}^{y}+\widehat{\varepsilon}_{i}^{z}+\frac{N}{K_{i}}\right).$
Consequently, using (108), the time-discretization contribution is of order $O(h^{1/2})$ under Assumption 1. Under Assumption 2, the $Y$-component has order $O(h)$, while the $Z$-component remains $O(h^{1/2})$; hence the combined estimate in (109) is still of order $O(h^{1/2})$ unless a sharper estimate for $\varepsilon_{Z}(h)$ is available.
Proof.
From [3], we have
$\left\{\begin{aligned} \max_{0\leq i\leq N-1}\mathbb{E}\left[\left|Y_{t_{i}}-Y_{i}^{\pi,m}\right|^{2}\right]&=\varepsilon_{Y}(h),\\ \mathbb{E}\left[\sum_{i=0}^{N-1}\int_{t_{i}}^{t_{i+1}}\left|Z_{t}-Z_{i}^{\pi,m}\right|^{2}dt\right]&=\varepsilon_{Z}(h).\end{aligned}\right$. (110)
Moreover,
$\varepsilon_{Y}(h)=O(h^{1/2})\quad\text{under Assumption 1},\qquad\varepsilon_{Y}(h)=O(h)\quad\text{under Assumption 2},$
and
$\varepsilon_{Z}(h)=O(h^{1/2}).$
From (98) and (103), we have, for $i\in\{0,1,\ldots,N-1\}$,
$\overline{Y}_{i}^{\pi,m}-\mathcal{U}_{i}^{m}=$ $\ \mathbb{E}_{i}\left[Y_{i+1}^{\pi,m}-\widehat{\mathcal{Y}}_{i+1}^{m}\right]$ (111)
$+h\left(f(t_{i},X_{i}^{\pi,m},\overline{Y}_{i}^{\pi,m},Z_{i}^{\pi,m})-f(t_{i},X_{i}^{\pi,m},\mathcal{U}_{i}^{m},\widehat{\mathcal{V}}_{i}^{m})\right).$
Proceeding similarly as Step 1 in the proof of Theorem 1, by Young inequality, Cauchy-Schwarz inequality, and the Lipschitz condition with respect to $f$, we obtain
$\mathbb{E}\left[\left|\overline{Y}_{i}^{\pi,m}-\mathcal{U}_{i}^{m}\right|^{2}\right]$ $\leq(1+\gamma h)\mathbb{E}\left[\left|\mathbb{E}_{i}\left[Y_{i+1}^{\pi,m}-\widehat{\mathcal{Y}}_{i+1}^{m}\right]\right|^{2}\right]$ (112)
$\quad+2L^{2}h^{2}\left(1+\frac{1}{\gamma h}\right)\mathbb{E}\left[\left|\overline{Y}_{i}^{\pi,m}-\mathcal{U}_{i}^{m}\right|^{2}+\left|Z_{i}^{\pi,m}-\widehat{\mathcal{V}}_{i}^{m}\right|^{2}\right].$
@@
$\mathbb{E}\left[\left|Z_{i}^{\pi,m}-\widehat{\mathcal{V}}_{i}^{m}\right|^{2}\right]$ $\leq\frac{2d}{h}\mathbb{E}\left[\left|Y_{i+1}^{\pi,m}-\widehat{\mathcal{Y}}_{i+1}^{m}\right|^{2}\right]$ (113)
$\quad-\frac{2d}{h}\mathbb{E}\left[\left|\mathbb{E}_{i}\left[Y_{i+1}^{\pi,m}-\widehat{\mathcal{Y}}_{i+1}^{m}\right]\right|^{2}\right].$
Plugging (113) into (112) and choosing $\gamma=4dL^{2}$, we obtain, for $h$ small enough,
$\mathbb{E}\left[\left|\overline{Y}_{i}^{\pi,m}-\mathcal{U}_{i}^{m}\right|^{2}\right]\leq(1+Ch)\mathbb{E}\left[\left|Y_{i+1}^{\pi,m}-\widehat{\mathcal{Y}}_{i+1}^{m}\right|^{2}\right].$ (114)
Next, by Young inequality as in Step 1 of the proof of Theorem 1, we have
$\mathbb{E}\left[\left|\overline{Y}_{i}^{\pi,m}-\overline{\mathcal{Y}}_{i}\right|^{2}\right]$ $\leq(1+Ch)\mathbb{E}\left[\left|Y_{i+1}^{\pi,m}-\widehat{\mathcal{Y}}_{i+1}^{m}\right|^{2}\right]$ (115)
$\quad+CN\mathbb{E}\left[\left|\overline{\mathcal{Y}}_{i}-\mathcal{U}_{i}^{m}\right|^{2}\right].$
@@
By the same arguments as in Step 2 in the proof of Theorem 1, it follows that, for $h$ small enough,
@@
$\mathbb{E}\left[\left|\overline{\mathcal{Y}}_{i}-\mathcal{U}_{i}^{m}\right|^{2}\right]\leq C\left(\widehat{\varepsilon}_{i}^{y}+h\widehat{\varepsilon}_{i}^{z}+\frac{1}{K_{i}}\right).$ (116)
Since
$Y_{i}^{\pi,m}=\max\left\{\overline{Y}_{i}^{\pi,m},\Phi(X_{i}^{\pi,m})\right\},\qquad\widehat{\mathcal{Y}}_{i}=\max\left\{\overline{\mathcal{Y}}_{i},\Phi(X_{i}^{\pi,m})\right\},$
and
$|\max(a,c)-\max(b,c)|\leq|a-b|,$
we obtain from (115) and (116) that
$\mathbb{E}\left[\left|Y_{i}^{\pi,m}-\widehat{\mathcal{Y}}_{i}\right|^{2}\right]$ $\leq(1+Ch)\mathbb{E}\left[\left|Y_{i+1}^{\pi,m}-\widehat{\mathcal{Y}}_{i+1}^{m}\right|^{2}\right]$ (117)
$\quad+C\left(\widehat{\varepsilon}_{i}^{y}+h\widehat{\varepsilon}_{i}^{z}+\frac{1}{K_{i}}\right).$
Using the terminal identity
$Y_{N}^{\pi,m}=g(X_{N}^{\pi,m})=\widehat{\mathcal{Y}}_{N},$
and applying the discrete Gronwall inequality backward in time, we derive
$\max_{0\leq i\leq N-1}\mathbb{E}\left[\left|Y_{i}^{\pi,m}-\widehat{\mathcal{Y}}_{i}\right|^{2}\right]\leq C\sum_{i=0}^{N-1}\left(\widehat{\varepsilon}_{i}^{y}+h\widehat{\varepsilon}_{i}^{z}+\frac{1}{K_{i}}\right).$ (118)
Combining (118) with (110), we deduce
$\max_{0\leq i\leq N-1}\mathbb{E}\left[\left|Y_{t_{i}}-\widehat{\mathcal{Y}}_{i}\right|^{2}\right]\leq C\varepsilon_{Y}(h)+C\sum_{i=0}^{N-1}\left(\widehat{\varepsilon}_{i}^{y}+h\widehat{\varepsilon}_{i}^{z}+\frac{1}{K_{i}}\right).$ (119)
This proves the estimate for the $Y$-component.
It remains to estimate the $Z$-component. By Young inequality and the definition of $\varepsilon_{Z}(h)$, we have
$\mathbb{E}\left[\sum_{i=0}^{N-1}\int_{t_{i}}^{t_{i+1}}\left|Z_{s}-\widehat{\mathcal{Z}}_{i}\right|^{2}ds\right]$ (120)
$\leq 2\mathbb{E}\left[\sum_{i=0}^{N-1}\int_{t_{i}}^{t_{i+1}}\left|Z_{s}-Z_{i}^{\pi,m}\right|^{2}ds\right]+2\sum_{i=0}^{N-1}h\mathbb{E}\left[\left|Z_{i}^{\pi,m}-\widehat{\mathcal{Z}}_{i}\right|^{2}\right]$
$\leq 2\varepsilon_{Z}(h)+2\sum_{i=0}^{N-1}h\mathbb{E}\left[\left|Z_{i}^{\pi,m}-\widehat{\mathcal{Z}}_{i}\right|^{2}\right].$
For each $i=0,\ldots,N-1$, since
@@
$Z_{i}^{\pi,m}=\mathbb{E}_{i}\left[Y_{i+1}^{\pi,m}\frac{(\Delta W_{i}^{m})^{\top}}{h}\right],\qquad\widehat{\mathcal{V}}_{i}^{m}=\mathbb{E}_{i}\left[\widehat{\mathcal{Y}}_{i+1}^{m}\frac{(\Delta W_{i}^{m})^{\top}}{h}\right],$
Cauchy-Schwarz inequality yields
$h\mathbb{E}\left[\left|Z_{i}^{\pi,m}-\widehat{\mathcal{V}}_{i}^{m}\right|^{2}\right]\leq C\mathbb{E}\left[\left|Y_{i+1}^{\pi,m}-\widehat{\mathcal{Y}}_{i+1}^{m}\right|^{2}\right].$ (121)
Moreover, by the same regression argument as in Step 2 of the proof of Theorem 1, together with the Monte Carlo estimate for the $Z$-label, we have
$h\mathbb{E}\left[\left|\widehat{\mathcal{V}}_{i}^{m}-\widehat{\mathcal{Z}}_{i}\right|^{2}\right]\leq C\left(h\widehat{\varepsilon}_{i}^{z}+\frac{1}{K_{i}}\right).$ (122)
Combining (121) and (122), we obtain
$h\mathbb{E}\left[\left|Z_{i}^{\pi,m}-\widehat{\mathcal{Z}}_{i}\right|^{2}\right]\leq C\mathbb{E}\left[\left|Y_{i+1}^{\pi,m}-\widehat{\mathcal{Y}}_{i+1}^{m}\right|^{2}\right]+C\left(h\widehat{\varepsilon}_{i}^{z}+\frac{1}{K_{i}}\right).$ (123)
Summing (123) over $i=0,\ldots,N-1$, and using (118), we get
$\sum_{i=0}^{N-1}h\mathbb{E}\left[\left|Z_{i}^{\pi,m}-\widehat{\mathcal{Z}}_{i}\right|^{2}\right]$ $\leq C\sum_{i=0}^{N-1}\mathbb{E}\left[\left|Y_{i+1}^{\pi,m}-\widehat{\mathcal{Y}}_{i+1}^{m}\right|^{2}\right]+C\sum_{i=0}^{N-1}\left(h\widehat{\varepsilon}_{i}^{z}+\frac{1}{K_{i}}\right)$ (124)
$\leq C\sum_{i=0}^{N-1}\left(N\widehat{\varepsilon}_{i}^{y}+\widehat{\varepsilon}_{i}^{z}+\frac{N}{K_{i}}\right).$
Inserting (124) into (120), we obtain
$\mathbb{E}\left[\sum_{i=0}^{N-1}\int_{t_{i}}^{t_{i+1}}\left|Z_{s}-\widehat{\mathcal{Z}}_{i}\right|^{2}ds\right]\leq C\varepsilon_{Z}(h)+C\sum_{i=0}^{N-1}\left(N\widehat{\varepsilon}_{i}^{y}+\widehat{\varepsilon}_{i}^{z}+\frac{N}{K_{i}}\right).$ (125)
Finally, adding (119) and (125), and using
$\sum_{i=0}^{N-1}\left(\widehat{\varepsilon}_{i}^{y}+h\widehat{\varepsilon}_{i}^{z}+\frac{1}{K_{i}}\right)\leq C\sum_{i=0}^{N-1}\left(N\widehat{\varepsilon}_{i}^{y}+\widehat{\varepsilon}_{i}^{z}+\frac{N}{K_{i}}\right),$
we obtain (109). This completes the proof. ∎
@@ 5 Numerical experiments
In this section, we demonstrate the performance of Algorithm 1 to solve the high dimensional nonlinear PDEs (1) via two numerical examples.
@@
For notational clarity in the numerical tables, $u(0,x_{0})$ denotes the exact analytical value at the initial point. For the $\ell$-th independent run, we denote the corresponding numerical estimate by $u_{\ell}^{\pi}(0,x_{0})$. The sample mean reported in the tables is denoted by
$\overline{u}^{\,\pi}(0,x_{0}):=\frac{1}{N_{\mathrm{rep}}}\sum_{\ell=1}^{N_{\mathrm{rep}}}u_{\ell}^{\pi}(0,x_{0}),\qquad N_{\mathrm{rep}}=10.$
The reported mean absolute error and relative error are computed by
$\mathbb{E}\left[\left|u(0,x_{0})-u^{\pi}(0,x_{0})\right|\right]\approx\frac{1}{N_{\mathrm{rep}}}\sum_{\ell=1}^{N_{\mathrm{rep}}}\left|u(0,x_{0})-u_{\ell}^{\pi}(0,x_{0})\right|,$
and
$\mathrm{Rel.\ err.}=\frac{\left|u(0,x_{0})-\overline{u}^{\,\pi}(0,x_{0})\right|}{\left|u(0,x_{0})\right|},$
respectively.
@@
$\begin{cases}\displaystyle\frac{\partial u}{\partial t}(t,x)+\frac{\sigma^{2}}{2}\sum_{\ell=1}^{d}D^{2}_{x_{\ell}}u(t,x)+\Big(u(t,x)-\frac{d+2}{2d}\Big)\Big(\sigma\sum_{\ell=1}^{d}D_{x_{\ell}}u(t,x)\Big)=0,\\ \displaystyle u(T,x)=\frac{\exp\!\left(T+\tfrac{1}{d}\sum_{\ell=1}^{d}x_{\ell}\right)}{1+\exp\!\left(T+\tfrac{1}{d}\sum_{\ell=1}^{d}x_{\ell}\right)}.\end{cases}$ (126)
@@
In the numerical experiments we take the terminal time $T=1$, the number of time steps $N=10$, the Monte Carlo samples $M=10000$, and the initial state $x_{0}=(0,0,\dots,0)^{\top}\in\mathbb{R}^{d}$. Thus, the exact solution of (126) at the initial point is $u(0,x_{0})=0.5$. The numerical estimate obtained in each independent run is denoted by $u_{\ell}^{\pi}(0,x_{0})$, and its sample mean is denoted by $\overline{u}^{\,\pi}(0,x_{0})$. The numerical solver is implemented using a deep learning framework where two sub-networks are trained at each time step to approximate the solution $u$ and its gradient $\nabla u$. Each network consists of two hidden layers with $d+110$ units and $\tanh$ activations. We utilize the Adam optimizer with a learning rate of $5\times 10^{-4}$ and conduct 6,000 training iterations per step to ensure the convergence of the loss function.
(a) $u(t,x)$ and its estimate at time $t=0.2$.
(b) $u(t,x)$ and its estimate at time $t=0.4$.
(c) $u(t,x)$ and its estimate at time $t=0.6$.
(d) $u(t,x)$ and its estimate at time $t=0.8$.
Fig. 1: Estimated solution $\overline{\mathfrak{u}}(t,x)$ obtained by DBR versus exact solution $u(t,x)$ for Example 1 with $d=1$.
@@
Scheme $d$ $u(0,x_{0})$ $\overline{u}^{\,\pi}(0,x_{0})$ Std. dev. $\mathbb{E}[|u-u^{\pi}|]$ Rel. err.
@@
Table 1 complements these visual findings by providing a quantitative comparison between the proposed DBR algorithm and the existing DBDP1 method in [29] across multiple high dimensions ($50,80,100$ and $200$). For each scheme and dimension, we report the exact value $u(0,x_{0})$, the sample mean $\overline{u}^{\,\pi}(0,x_{0})$ over 10 independent runs, the empirical standard deviation $\mathrm{Std.\ dev.}(u^{\pi}(0,x_{0}))$, the mean absolute error
$\mathbb{E}\left[\left|u(0,x_{0})-u^{\pi}(0,x_{0})\right|\right],$
and the relative error
$\mathrm{Rel.\ err.}=\frac{\left|u(0,x_{0})-\overline{u}^{\,\pi}(0,x_{0})\right|}{\left|u(0,x_{0})\right|}.$
Regarding accuracy, the DBR algorithm maintains a relative error consistently below $4\%$ across all tested dimensions, significantly outperforming the DBDP1 method which exhibits errors as high as $13.9\%$ at dimension $50$ and $9.7\%$ at dimension $80$. More importantly, the mean absolute error, which represents the average magnitude of the absolute deviation between the true solution and the numerical approximation, is consistently smaller for the DBR method across all dimensions. For instance, at dimension $100$, DBR achieves a mean absolute error of $0.008006$ compared to DBDP1’s $0.045578$, indicating that the DBR algorithm produces approximations that are substantially closer to the true solution on average. This observation is consistent with the motivation of the DBR scheme, namely that conditional-expectation-based labels can reduce pathwise noise in the training targets.
Example 2. (see [17]) Consider the following high-dimensional nonlinear PDE with an unbounded solution and a more complex structure:
@@
with $\hat{k}(t,x)=\partial_{t}u+\frac{1}{2d}\operatorname{Tr}\left(D_{x}^{2}u\right)+\frac{u}{\sqrt{d}}D_{x_{i}}u+\frac{u^{2}}{2}$. Thus, the solution of PDE (127) is given by $u(t,x)=\frac{T-t}{d}\sum_{i=1}^{d}\left(\sin\left(x_{i}\right)1_{x_{i}<0}+x_{i}1_{x_{i}\geq 0}\right)+\cos\left(\sum_{i=1}^{d}ix_{i}\right).$
Example 2 presents a more challenging class of high-dimensional nonlinear PDEs characterized by an unbounded solution and a complex structural form. We note that Example 2 involves an unbounded solution. Hence this benchmark is used to test the robustness of the proposed algorithm beyond the bounded classical setting. The theoretical estimates above are stated under the regularity and integrability conditions specified in Sections 2 and 3.
To comprehensively evaluate the performance of the proposed DBR algorithm, we conduct systematic numerical experiments across dimensions $d=1,2,8,15,20$ with varying time steps $N$, where the initial point is fixed as $x_{0}=0.5\mathbf{1}_{d}$ for all cases. The neural network architecture consists of two hidden layers with hyperbolic tangent activation functions, where each hidden layer contains $d+10$ units. The model is trained using a batch size of 400 and a total of 3000 iterations.
(a) $u(t,x)$ and its estimate at time $t=0.3$.
(b) $u(t,x)$ and its estimate at time $t=0.5$.
(c) $u(t,x)$ and its estimate at time $t=0.7$.
(d) $u(t,x)$ and its estimate at time $t=0.9$.
Fig. 2: Estimated solution $\overline{\mathfrak{u}}(t,x)$ obtained by DBR versus exact solution $u(t,x)$ for Example 2 with $d=1$.
@@
Scheme NN $u(0,x_{0})$ $\overline{u}^{\,\pi}(0,x_{0})$ Std. dev. $\mathbb{E}[|u-u^{\pi}|]$ Rel. err.
@@
Scheme NN $u(0,x_{0})$ $\overline{u}^{\,\pi}(0,x_{0})$ Std. dev. $\mathbb{E}[|u-u^{\pi}|]$ Rel. err.
@@
Scheme NN $u(0,x_{0})$ $\overline{u}^{\,\pi}(0,x_{0})$ Std. dev. $\mathbb{E}[|u-u^{\pi}|]$ Rel. err.
@@
Scheme NN $u(0,x_{0})$ $\overline{u}^{\,\pi}(0,x_{0})$ Std. dev. $\mathbb{E}[|u-u^{\pi}|]$ Rel. err.
@@
Scheme NN $u(0,x_{0})$ $\overline{u}^{\,\pi}(0,x_{0})$ Std. dev. $\mathbb{E}[|u-u^{\pi}|]$ Rel. err.